SuperDuper

An Amiga Disk Copier

Version 3.0

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991,1992,1993 Sebastiano Vigna

by Sebastiano Vigna

1 Introduction

SuperDuper is a disk copier/formatter that tries to be to disk handling what Mostra is to IFF displaying: a fast, compact, system-friendly tool which combines speed, features, and some bells and whistles to make your life easier.

By fast I mean exactly what you’re hoping�blazingly fast. A disk is usually copied and veri�ed in less than 100s. Without verify, the time drops to 69s. You can buffer a disk in ram in less than 36s, and then making a veri�ed copy takes 67s, while a non-veri�ed copy takes less than 36s. Adding another destination drive increases veri�ed copy times by 34s, but hardly changes non-veri�ed copy times (the Amiga can write more than one drive at a time; I just need a few tenths of a second in order to measure the drive speed and step the heads). Thus, if you really trust your drives and your media you can make four copies in 38s. These timings can vary with the system con�guration, the multitasking overhead, the disposition of the blocks on the surface of the disk, the state of the Date option (which requires a separate write on the root block track for each disk), the dma access of the custom chips and the display features enabled.

1.1 Changes

SuperDuper 3.0 is in many ways a completely new program. This is why I strongly suggest to all previous users to entirely reread the documentation.

Now the program works only under release 2.04 and beyond of the operating system. Minor maintenance releases of SuperDuper 2.0x will probably be distributed for the users who do not still have Release 2, but you should assume that no programming effort from my part will be ever spent for writing something working under 1.3. Moreover, some features of 3.0 are used if present (for instance, scalable checkmarks and radio buttons gadgets).

The main new features are high density drive support, xpk compression support (for more details, see Section 4.2 [The Buffering System], page 10), a complete graphical display of the copy process status, and much more �exible buffering.

1.2 Main Features

• SuperDuper copies, formats and checks from/to any combination of Amiga drive(s).

• SuperDuper can buffer a disk in ram, allowing for any number of duplications while reading the source disk only once. The combination of destination drives can be changed at each pass. If you have a hard disk, you can create on it an IFF image �le that will act as a buffer. This �le can be saved and reused many times. Also, all kinds of virtual disks are supported for buffering (VD0:, RAD:, FMS: ... ).

• SuperDuper checksums the ram buffer. If some badly written program is trashing your memory, you are alerted. Thus, buffered copies are as safe as direct copies.

• SuperDuper also checks its internal dma buffers at each write.

• SuperDuper can use any xpk library in order to compresss the information it stores in ram. You can choose your preferred compression algorithm, depending on the speed of your machine and on the available memory. With most compressors, write times are (almost) unaffected.

• SuperDuper will automatically retry tracks which produce a verify error. The number of retries is programmable. A complete graphical rendition of the status of retries and errors is given in the Info Window, and on request detailed error information printing is available.

• SuperDuper is highly system-friendly�the use of cpu time is negligible, so you can multitask ef�ciently.

• SuperDuper has the option of incrementing the creation date of the copy so AmigaDOS doesn’t get confused. If, however, the option is switched on and the disk is not an AmigaDOS disk, SuperDuper won’t increment the date.

• SuperDuper is faster than diskcopy�actually it pushes the drives to their limits. At the time of this writing, SuperDuper is the fastest Amiga copier both from a �pure� (physical time) and from a �per-copy� (real time for each copy when a big number of copies of the same disk is produced) point of view.

• SuperDuper alerts the user with sound (and optionally voice) about the operations in progress�so you can really be doing something else!

• SuperDuper can format both OFS, FFS and DCFS disks.

• SuperDuper displays a list of the last disks copied. If you do a lot of copying, you’ll �nd this feature more than a little useful.

• SuperDuper can manage the Amiga drives without help from the trackdisk.device. Through the supplied utility SDBootInstall, you can create a boot disk which will keep the system away from your drives, giving you back more than 30K